# MimicTouch: Leveraging Multi-modal Human Tactile Demonstrations for Contact-rich Manipulation

### v3

@@ 5 Limitation and Future Work
MimicTouch pioneers the pathway to learning human tactile-guided control strategies from human tactile demonstrations. However, it still has several limitations for future improvements. Firstly, MimicTouch still requires several hours to RL fine-tune the policy to address the embodiment gap and enrich the contact reasoning. We will explore a better representation learning method to further reduce the gap between humans and robots. Secondly, zero-shot generalization on soft objects can potentially be enhanced by collecting a small amount of new human tactile demonstrations for fine-tuning the offline policy. Thirdly, MimicTouch is task-specific and can not directly generalize human’s tactile-guided control strategy to other contact-rich manipulation tasks. One potential solution is to learn a generalizable tactile-based dynamic model, rather than a task-specific control policy. Finally, the method of learning to perform contact-rich manipulation tasks from human tactile demonstrations could be extended to other robot tasks, such as dexterous manipulation, bimanual manipulation, and deformable object manipulation.
@@ 6 Conclusion
We presented MimicTouch, a multi-modal imitation learning framework that: i) enables humans to perform tactile demonstrations with their hands without a robot in the loop ii) learns from such demonstrations and safety transfer to robot with non-parametric imitation learning, and iii) improves the policy performance with residual-based online RL to bridge the human-robot embodiment gap. We show that MimicTouch enables high-throughput data collection and achieves high success rate and generalization across a wide range of two-piece insertion and assembly tasks.
@@ Appendix D Details of RL training
2:Input: step size sequences $\{\beta_{t}\}$, number of iterations $K$, Replay Buffer $D$
@@
6:     Obtain current trajectory $C_{k}$ by executing $\pi_{o}+\pi_{r}$
7:     To collect more data in $D$: $D\leftarrow D\cup C_{k}$
8:     Combine $D_{k}$ and $C_{k}$ to form batch $B_{k}$ for update
9:     for all $(s,a_{i},r,s^{\prime})\in B_{k}$ do
10:         $a_{o}\leftarrow\pi_{o}(s)$triangleright Obtain offline action
11:         $a_{r}\leftarrow\pi_{r}(s,a_{o})$triangleright Obtain residual action
@@
13:         $Q_{r}\leftarrow Q_{r}(s,a_{o})+r+\gamma Q_{r}(s^{\prime},\pi_{o}(s^{\prime}))$
14:         $\pi_{r}\leftarrow\pi_{r}-\alpha\nabla_{\pi_{r}}L(\pi_{r})$triangleright Update residual policy with gradient step

### v4

@@ 5 Limitation and Future Work
MimicTouch pioneers the pathway to learning human tactile-guided control strategies from human tactile demonstrations. However, it still has several limitations for future improvements. Firstly, MimicTouch still requires several hours to refine the policy for addressing the embodiment gap. We will explore a better representation learning method to reduce the gap between humans and robots. Secondly, MimicTouch is task-specific and can not directly generalize human’s tactile-guided control strategy to other tasks. One potential solution is to learn a generalizable tactile-based dynamic model for different tasks. Thirdly, the method of learning to perform contact-rich manipulation tasks from human tactile demonstrations could be extended to other robot tasks, such as dexterous manipulation, bimanual manipulation, and soft object manipulation.
@@ 6 Conclusion
We presented MimicTouch, a multi-modal imitation learning framework that: i) enables humans to perform tactile demonstrations with their hands without a robot in the loop, ii) learns from such demonstrations and safely transfers to robot with non-parametric imitation learning, and iii) improves the policy performance with residual-based online RL to bridge the human-robot embodiment gap. We show that MimicTouch enables high-throughput data collection and achieves high success rate and generalization across a wide range of two-piece insertion and assembly tasks.
@@ Appendix D Details of RL training
2:Input: Step size sequence $\{\beta_{t}\}$, number of iterations $K$, replay buffer $D$
@@
6:     Execute combined policy $\pi_{o}+\pi_{r}$ to collect trajectory $C_{k}=(s,\hat{a},a_{o},r_{t},s^{\prime})$
7:     Add new data to replay buffer: $D\leftarrow D\cup C_{k}$
8:     Form combined batch $B_{k}=D_{k}\cup C_{k}$ for updates
9:     for all $(s,\hat{a},a_{o},r_{t},s^{\prime})\in B_{k}$ do
10:         $a_{o}\leftarrow\pi_{o}(s)$triangleright Compute offline action
11:         $a_{r}\leftarrow\pi_{r}(s,a_{o})$triangleright Compute residual action
@@
13:         Update residual Q-function:
$Q_{r}\leftarrow Q_{r}(s,\hat{a},a_{o})+\alpha\big{(}r_{t}+\gamma\max_{\hat{a}^{\prime}}Q_{r}^{\prime}(s^{\prime},\hat{a}^{\prime},a_{o}^{\prime})-Q_{r}(s,\hat{a},a_{o})\big{)}$
14:         Update residual policy $\pi_{r}$:
$\pi_{r}\leftarrow\pi_{r}-\alpha\nabla_{\pi_{r}}L(\pi_{r}),$
where
$L(\pi_{r})=-\mathbb{E}_{(s,\hat{a},a_{o},r_{t},s^{\prime})\sim B_{k}}[Q_{r}(s,\hat{a},a_{o})]-\alpha\log\pi(a|s)$